OpenStudy (ria23):

Find the intersection points, identify the type of equation, and the possible number of solutions. x^2+y^2-16y+39=0 y^2-x^2-9=0 I'm stuck... ;-;

OpenStudy (anonymous):

yuck i guess we can add them maybe to start?

OpenStudy (ria23):

xD Haha, sure, why not. c:

OpenStudy (anonymous):

i get \[2y^2-16y+25=0\]

OpenStudy (ria23):

For... The first one, right?

OpenStudy (anonymous):

i just added in a column

OpenStudy (ria23):

Or combined?

OpenStudy (anonymous):

the \(x^2\) terms add to zero, leaving what i wrote

OpenStudy (anonymous):

ok that is all wrong, lets try something else

OpenStudy (ria23):

Would solving by substitution or addition/elimination work? .-.

OpenStudy (anonymous):

i was trying to use elimination but i got the wrong answers for \(y\) for some reason now sure why it did not work, maybe i added incorrectly

OpenStudy (anonymous):

oh because i am an idiot!

OpenStudy (ria23):

xD Lies.

OpenStudy (anonymous):

i wrote \[2y^2-16y+25=0\] when it should have been \[2y^2-16y+30=0\]

OpenStudy (anonymous):

first of before we go any further, do you see how i got that equation ?

OpenStudy (ria23):

Yhu, added down the column again?

OpenStudy (anonymous):

right only this time i did it correctly now we have an easy quadratic equation to solve \[2y^2-16y+30=0\] or even easier \[y^2-8y=15=0\] care to solve it? it factors

OpenStudy (ria23):

Should that be +15?

OpenStudy (anonymous):

yeah i must be tired \[y^2-8y+15=0\]

OpenStudy (ria23):

Haha would it be \[(y+3)^{2}(y-5y)^{2}\]?

OpenStudy (anonymous):

?? i thought i was tired no it would be \((y-3)(y-5)=0\)

OpenStudy (ria23):

; - ; Can yhu explain why?

OpenStudy (anonymous):

you mean can i explain how to factor? no, you just have to grind it til you find it but when you have \[y^2-8y+15=0\] you are looking for two numbers whose product is \(15\) and whose sum is \(-8\) and \(-3,-5\) fit that bill nicely

OpenStudy (ria23):

-.- I switched the signs. :c Thank yhu... So after we factor it, would yhu solve for y?

OpenStudy (anonymous):

yeah and get \(y=3,y=5\) pretty much in your head now we have to go back and find the x values for both there will be four solutions all together, two for \(y=3\) and two for \(y=5\)

OpenStudy (ria23):

OS froze. ;c So we would have to solve it 4 times?

OpenStudy (anonymous):

lets use this one \[y^2-x^2-9=\] with \(y =3\) you get \[9-x^2-9=0\\ x^2=0\\ x=0\] so \((0,3)\) is one of them

OpenStudy (anonymous):

no you don't have to use both equations, only one

OpenStudy (anonymous):

but you have to solve it when \(y=3\) and \(y=5\) \(y=3\) only gives on solution, not two, because \(x^2=0\) only has one solution now lets try with \(y=5\)

OpenStudy (anonymous):

\[y^2-x^2-9=0\\ 25-x^2-9=0\\ x^2=16\\ x=\pm4\]

OpenStudy (anonymous):

that wasn't too bad

OpenStudy (ria23):

And then we'd plug in the values we just got for x to solve for y? c:

OpenStudy (anonymous):

lol no dear we have already solved for y it is \(y=3\) or \(y=5\)

OpenStudy (ria23):

Maybe I'm tired too. >~<

OpenStudy (anonymous):

we did that first then if \(y=3\) we found that \(x=0\) so one solution is \((0,3)\)

OpenStudy (ria23):

So we come out with 3 solutions...

OpenStudy (anonymous):

yeah me too but to finish it we found it \(y=5\) then \(x=4\) or \(x=-4\) making two more solutions:\[(4,5),(-4,5)\]

OpenStudy (ria23):

Total.

OpenStudy (anonymous):

yup 3 total

OpenStudy (ria23):

And one last part for the question, Which type of equations are these? If I say my guess yhu can't laugh, I don't know if it's right. ; 3; Would it be a system of equations?

OpenStudy (anonymous):

lololololoooollll etc etc

OpenStudy (anonymous):

no

OpenStudy (anonymous):

one is a circle, the other a hyperbola

OpenStudy (anonymous):

\[x^2+y^2-16y+39=0\] is a circle \[y^2-x^2-9=0\] is a hyperbola

OpenStudy (ria23):

I have one I did earlier... Would yhu mind telling me which is which? I got the answersm it was a little easier than this, but I don't know what they are... \[\frac{ x^{2} }{ 4 }+y^{2}=1\]\[y=x+1\] I was thinking maybe hyperbola... and line?

OpenStudy (anonymous):

a line for sure

OpenStudy (anonymous):

the first is not a hyperbola as it is \[\frac{x^2}{4}\color{red}+y^2=1\]

OpenStudy (anonymous):

it is an ellipse if it had a minus sign there it would have been a hyperbola

OpenStudy (ria23):

Oh... I get those two mixed a lot. Haha, well thank yhu for everything! n.n Yhur awesome. And yhu should probably get some sleep. Haha

OpenStudy (anonymous):

yw gnight you too

OpenStudy (ria23):

'Night c: